# GUE correlators and large genus asymptotics

### v1

@@ 1. Introduction
In this paper, using the matrix-resolvent formula for the GUE (Gaussian unitary ensemble) correlators in [5] (which are closely related to enumerations of ribbon graphs), we will study certain large genus asymptotic behaviors and certain rationality properties. Our method is similar to [7]. In particular, we will obtain large genus asymptotics of enumerations of ordinary graphs and ribbon graphs with 1 face.
@@
(1) $\langle\text{tr}M^{i_{1}}\cdots\text{tr}M^{i_{n}}\rangle_{c}(N):=\frac{\int_{\mathcal{H}(N)}\text{tr}M^{i_{1}}\cdots\text{tr}M^{i_{n}}e^{-\frac{1}{2}\text{tr}M^{2}}\mathop{}\!\mathrm{d}M}{\int_{\mathcal{H}(N)}e^{-\frac{1}{2}\text{tr}M^{2}}\mathop{}\!\mathrm{d}M},$

### v2

@@ 1. Introduction
In this paper, using the matrix-resolvent formula [8] for the GUE (Gaussian unitary ensemble) correlators (which are closely related to enumerations of ribbon graphs), we will study certain large genus asymptotic behaviors and certain rationality properties. Our method is similar to [10]. In particular, we will obtain large genus asymptotics of enumerations of ordinary graphs and ribbon graphs with 1 face. We note that Pasquetti and Schiappa [14] analyzed large genus behavior of certain GUE free energy from the perspective of resurgence theory; see also [1] for a comprehensive review on resurgence. We also note that a deep result on enumeration of triangulations in large genus was obtained in [6].
@@
(1) $\langle\text{tr}M^{i_{1}}\cdots\text{tr}M^{i_{n}}\rangle(N):=\frac{\int_{\mathcal{H}(N)}\text{tr}M^{i_{1}}\cdots\text{tr}M^{i_{n}}e^{-\frac{1}{2}\text{tr}M^{2}}\mathop{}\!\mathrm{d}M}{\int_{\mathcal{H}(N)}e^{-\frac{1}{2}\text{tr}M^{2}}\mathop{}\!\mathrm{d}M},$
@@
The connected GUE correlators $\langle\text{tr}M^{i_{1}}\cdots\text{tr}M^{i_{n}}\rangle_{c}(N)$ can be defined through the relation
(4) $\langle\text{tr}M^{i_{1}}\cdots\text{tr}M^{i_{n}}\rangle(N)=\sum_{\mathcal{P}\text{ partition of }\{1,\cdots,n\}}\prod_{I\in\mathcal{P}}\left\langle\prod_{j\in I}\mathrm{tr}M^{i_{j}}\right\rangle_{c}(N).$
@@ 3. Asymptotic Formula on Enumeration of Ribbon Graphs with 1 Face
We note that an upper bound for correlators in topological recursion was given in [5] in a general context. In our situation, the upper bound given in [5] involves a coefficient with an unspecified dependence on the genus $g$; on the other hand, our estimate here gives the precise leading asymptotics.